• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:01:11 on June19,2002

    Entry number 86429

    keyword=Shift Summary

    Day shift: 6/19/02
    John LeRose (SL)
    Ron Gilman (TO)
    Brian Milbraith
    Jing Yuan

    At start of shift doing production running at Kinematics Q2_f40r.
    Running at 80 microamps, current limited by a problem with the high power heaters in the target (see halog 86398)
    run 1643 in progress at start of shift.
    8:50 run 1644 started
    9:40 J-P declares there isn't any problem with the high power heater. Requested 100 microamps from MCC
    9:50 run 1645 started
    10:00 requested 110 microamps, without any heroics, got 107. We'll take it!
    10:35 run 1646 started (with header corrected for Q2_f40r)
    11:13 optics runs 1647, noticed beam spot was off center
    Had MCC recenter beam
    11:19 optics run 1648 with more centered beam
    11:24 switch to kinematics Q2_f50r
    Arun does his harp tests during switchover
    Right arm: 58.509 degrees P0=1.6870 GeV/c
    Left arm: 20.34 degrees P0=3.5829 GeV/c
    11:38 Arun is finished and we're ready to go
    11:48 Run 1649, optics run, started.
    12:00 putting in LD2 target
    12:08 Run 1650 started
    12:58 Run 1651 started
    13:57 Run 1652 started
    14:12 ROC14 crashed, reset and restarted coda
    14:16 Run 1653 started
    14:29 more problems with ROC14
    14:41 Run 1654
    14:55 MCC says They're having problems with injector steering. Low current for 5-10 minutes. was less than 5 minutes and not very low.
    14:57 More ROC14 problems
    Having trouble getting CODA back up
    Reset ROC's with help from Bob and Bodo. Looked like they did all the same things we did but somehow it works for them??? Suspect maybe we didn't wait long enough after resets. You must check and make sure the resets are complete before restarting CODA. Bodo likes to type "i" in the ROC control screens. You get a nice table if all is well. If you get an exception you probably want to reset again.
    15:20 run 1656 started
    Still running at end of shift